OpenStudy (anonymous):

Kim walked to the grocery store and back home.The grocery store is 2 5/8 miles from her house. The next day kim walked twice as far to the post office and then back home.How far did she walk in all?

OpenStudy (rosedewittbukater):

2 5/8 x 2 = 5 1/4 This is the distance she walked the first day to the grocery store and back. It is also the distance one way to the post office, since it is twice as far. 5 1/4 x 2 = 10. 1/2 This is the distance she walked the second day to the post office and back. Now add the distance she walked the first day with the distance she walked the second day. 10 1/2 + 5 1/4 = ?